Could someone help me with this one? I have trouble with word problems. You are greatly appreciated.
One year ago Walt invested $12,000. He invested part of the money at 7% and the rest at 9%. He made a total of $970 in interest. How much was invested at 7%?

Answer by rfer(16322)

let x=7% investment=$5500
let 12000-x=9% investment=$6500
.07x+.09(12000-x)=970
.07x+1080-.09x=970
-.02x=-110
x=5500